Soy bean 16% protein; corn meal 8% protein; How many pounds of each should be mixed together in order to get 320-lb mixture that is 11%protein.
Answer by nerdybill(7384)

Soy bean 16% protein; corn meal 8% protein; How many pounds of each should be mixed together in order to get 320-lb mixture that is 11% protein.
.
Let x = pounds of soy beans
and y = pounds of corn meal
.
Since we have two unknowns we MUST find two equations.
Equation 1 we extract from the fact that the total weight is 320 pounds:
x + y = 320
.
Equation 2 we extract from the amount of protein found in each:
.16x + .08y = .11(320)
.
Solve equation 1 for y:
x + y = 320
y = 320 - x
.
Substitute the above into equation 2 and solve for x:
.16x + .08y = .11(320)
.16x + .08(320 - x) = .11(320)
.16x + 25.6 - .08x = 35.2
.08x + 25.6 = 35.2
.08x = 9.6
x = 120 pounds (soy beans)
.
To find corn meal, plug the above into equation 1:
x + y = 320
120 + y = 320
y = 200 pounds (corn meal)